# The Population II Extragalactic Distance Scale: A TRGB Distance to the Fornax Cluster with JWST

### v1

@@ 1 Introduction
For at least a half a century, the Virgo Cluster has played an important role in the establishment of the extragalactic distance scale (Sandage & Tammann, 1976). However, the depth of the virializing region of this cluster, about ±$\pm 10\%$ of its distance, is an appreciated hindrance, made more difficult by the ongoing accretion of new members (Tully & Shaya, 1984; Cantiello et al., 2024). The Fornax Cluster has played a more limited role, but it is more compact with a depth about ±$\pm 5\%$ of its distance and little evidence of recent accretion. Fornax is only slightly more distant than Virgo; surface brightness fluctuation (SBF) measurements for 43 galaxies in Fornax and 91 galaxies in the Virgo vicinity using the Advanced Camera for Surveys (ACS) on Hubble Space Telescope (HST), imply a Virgo/Fornax distance ratio $d_{F}/d_{V}=1.21\pm 0.02$ (Blakeslee et al., 2009).
@@
Of equal interest to our program is the establishment of a robust zero-point calibration for the SBF methodology. The SBF technique with JWST holds great promise in measuring distances out into the Hubble flow (Blakeslee et al., 2021a; Cantiello & Blakeslee, 2023), allowing for an alternative to Type Ia supernovae as the final rung on the distance ladder (Riess et al., 2022). The old, gas-poor systems appropriate for SBF measurements do not contain significant numbers of Cepheids, requiring calibration, instead, with the TRGB. Aside from the peculiar galaxy Centaurus A and the obscured Maffei I, there are no giant elliptical galaxies within the 10 Mpc limit available for single orbit TRGB measurements with HST (a couple lie just beyond 10 Mpc in the Leo Group but already required multiple HST orbit observations for sufficient signal). As a consequence, TRGB distances have played only a minor role in establishing the SBF scale (Blakeslee et al., 2021b). With JWST, the very high point-source signal/noise (S/N) acquired with moderate exposures for galaxies within 20 Mpc provide an abundance of candidates to establish a TRGB-SBF linkage. Here we provide TRGB measurements to the three Fornax Cluster galaxies from our program and a first look at SBF measurements with JWST.
@@ 2 Data
The three galaxies presented in this work are the first of 14 targets of Cycle 2 JWST program GO-3055 (Tully et al., 2023a) to link the TRGB and SBF scales through simultaneous observations tailored to each method. The split fields of the JWST NIRCam detector (Gardner et al., 2023; Rieke et al., 2023) are ideal for the experiment. One field covers the high surface brightness core of a galaxy for the SBF measurement, while the other field extends to the galactic halo where the TRGB measurement can be made with minimal confusion from crowding. SBF can also be measured in the peripheral fields, providing a direct overlap with TRGB. The primary observations are made in the $F090W$ and $F150W$ filters. $F090W$ is optimal for TRGB and SBF because of the minimal variance of RGB magnitudes at lower metallicities (Anand et al., 2024). $F150W$ is also a good choice for SBF in the high metallicity domain of the galactic cores because the fluctuation signal is very bright and near its peak, making $F150W$ SBF measurable to large distances. Simultaneous long-wavelength observations are made in both $F277W$ and $F356W$, although we do not make use of these filters in this work. The integration times are set by the requirements of the TRGB observations. The SBF signal is abundantly strong in the same exposure time. The deep multi-wavelength imaging of the selected galaxies is also useful for many other ancillary science cases.
@@
2023-10-12 NGC 1399 o002_t002 F090W 8761.204
2023-10-12 NGC 1399 o002_t002 F150W 2104.408
2023-11-09 NGC 1380 o001_t001 F090W 8761.204
2023-11-09 NGC 1380 o001_t001 F150W 2104.408
2023-11-14 NGC 1404 o003_t003 F090W 8761.204
2023-11-14 NGC 1404 o003_t003 F150W 2104.408
Table 1: Observation log for the data used in this paper. All the images were processed with jwst_1140.pmap or jwst_1147.pmap context versions, which include the NIRCam flat-field and flux calibration updates introduced in jwst_1125.pmap and jwst_1126.pmap, respectively. The name of all observation IDs start with “jw03055-” and end with “_nircam_clear-f090w” or “_nircam_clear-f150w”.
Table 1 provides a summary of the data we use in this paper. All the data was reduced with context versions [^1] See https://jwst-crds.stsci.edu/display_all_contexts/ for more information on each version. 1140.pmap or 1147.pmap (with no meaningful changes for NIRCam imaging between the two), and contain all reference file updates as of January 2024, including the updates to the flat-fields and zeropoints provided in 1125.pmap and 1126.pmap, respectively.
Footprints of our NIRCam observations are shown in Fig. 1, overlaid on a one square degree image from the Digitized Sky Survey. From this figure, the extended spatial NIRCam coverage is clear. To further highlight the dynamic range probed by NIRCam in a single pointing, we show color ($F090W+F150W+F356W$) images of individual chips from the NGC 1380 visit in Fig. 2. The core of this galaxy is centered on the B4 chip; a small dust lane in the center is readily apparent. While individual stars cannot be resolved in the inner regions due to crowding, a few globular clusters are visible scattered throughout. Some background galaxies can also be seen. The two chips furthest from the core of the galaxy are A1 and A2, and we show a color image from the former. There are thousands of giant branch stars which populate this frame. Importantly, they can be easily separated from one another, due to the significantly lower surface brightness at this radius. A diverse population of background galaxies is also visible.
@@ 3 Data Reduction
For our photometry, we use the recommended parameters provided in the recent work of Weisz et al. (2024). We perform our PSF photometry directly on the Stage 2 cal images, while using a Stage 3 i2d image from the JWST pipeline as the “reference frame” for mutual image alignment. We choose $F090W$ for the reference frame, as our observations are the deepest in that filter. We filter our final object list based on the quality criteria presented in Anand et al. (2024), which are based on the recommendations provided by the ERS program team in Warfield et al. (2023). We only include sources which meet the following DOLPHOT criteria in both filters in our final object list (except for type, which is not a filter specific output): (1) Crowding $<$ 0.5; (2) $\mathrm{(Sharpness)^{2}}$leq 0.01; (3) Object Type $\leq$ 2; (4) S/N $\geq$ 5; (5) Error Flag $\leq$ 2. All of our PSF photometry is transformed to the Vega system (and not the Sirius–Vega variant).
We opt to utilize the “–etctime” option in the DOLPHOT/NIRCam module, which attempts to rectify the exposure times provided in the JWST image headers. The use of this option does not significantly affect the flux measurements themselves (due to the underlying image units of surface brightness); instead it aims to provide more realistic S/N estimates for each individual source, as the image headers appear to slightly overestimate the true on-sky exposure times (Weisz et al., 2024).
@@
Color-magnitude diagrams (CMDs) for each target can be seen in Fig. 3. These CMDs are limited to the two SW NIRCam chips that are furthest from the galaxy’s core (A1+A2). For NGC 1399, we further restrict the points to just those that are in the furthest quarter of each chip, for visual clarity. In each CMD, the vast majority of resolved stars are those on the red-giant branch, with the TRGB visible at $m_{F090W}\simeq 27$ mag ($D=19$ Mpc) for each target. The magnitude limit of the CMD for NGC 1399 is $\sim 0.3$ mag brighter than for the other two targets, likely due to crowding effects at the faint end (NGC 1399 is the cD galaxy in the Fornax Cluster, and thus more extended). It can also be seen that the red giant branch contains stars of a vast range of metallicities; for our TRGB work, we will only use the lower metallicity stars, which have little-to-no variation in their magnitude when observed in $F090W$ (see the discussion in §2 of Anand et al. 2024). A further description of the underlying stellar populations can be found in §4.6.
@@ 4 TRGB Measurements
For each target, we perform our measurements on each of the two chips that are furthest from the galaxy’s core. Due to the observational setup for our program, these are the A1 and A2 chips for all three targets. We analyze the data for both chips separately, reducing any concerns regarding the chip-to-chip zeropoint offsets. We note that for most standard stars, these offsets are at the level of 0.00–0.03 mag [^2] See the details provided at https://jwst-docs.stsci.edu/jwst-data-calibration-considerations/jwst-calibration-uncertainties for more information regarding calibration uncertainties.. For NGC 1399, we further limit our analysis to the least crowded quarter of each chip in an attempt to mitigate the effects of the (modest) crowding present (which is more of an issue for the edge-detection routine, as it does not employ the use of artificial stars).
@@
Figure 4: Color-magnitude diagrams and edge-detection measurements for each of our six fields in three galaxies. The maximum-likelihood TRGB results are also plotted for comparison, but are nearly indistinguishable from the edge-detection results except for the NGC 1399 fields (which suffer from a modest amount of photometric bias).
@@ 4.4 Distances
To provide an absolute scaling for our TRGB measurements, we utilize the JWST/NIRCam TRGB calibration provided by Anand et al. (2024). They use data taken in the outer regions of the megamaser host NGC 4258 to provide a direct calibration in our filter of interest ($F090W$). The baseline absolute calibration they provide is $M_{\mathrm{TRGB}}^{F090W}=-4.362\pm 0.033$ (stat) ±$\pm 0.045$ (sys) mag. They also note that this baseline calibration may be further tailored to individual measurement techniques: a value of $M_{\mathrm{TRGB}}^{F090W}=-4.347$ mag is more appropriate for techniques which employ the use of LF fitting with artificial stars (e.g. Makarov et al. 2006) and take into account photometric bias, errors, and incompleteness. We note that several modest changes in the recommended DOLPHOT parameters may result in minor offsets in the final photometry, though Weisz et al. (2024) note that the performance is comparable. [^4] An updated NIRCam TRGB calibration with the new DOLPHOT parameters and PSF models, along with a NIRISS TRGB calibration will be provided in a future, separate work (Anand et al. in prep). To account for any potential mismatches, we choose to include an additional 0.03 mag contribution to the overall systematic error term. The magnitude of this term is also provided by Anand et al. (2024), who obtain this value from several other works in the literature which examine the systematic effects of changing DOLPHOT parameters on the output photometry (Jang et al., 2021; Jang, 2023). Our adopted absolute calibration now becomes $M_{\mathrm{TRGB}}^{F090W}=-4.347\pm 0.033$ (stat) ±$\pm 0.054$ (sys) mag (see Table 3 in Anand et al. 2024 for all of the included error terms).
@@ 4.5 Comparison with Previous Distances
Figure 7: Comparison of two CMDs (down to S/N = 5) for NGC 1404 used to measure the TRGB distance to this galaxy. The left panel shows the reduction of the 34 orbit (crowding-limited) HST dataset, whereas the right panel shows the reduction of the 4.74 hour JWST dataset. The y-axes of the two panels are set to match at the measured levels of the TRGB (denoted with dashed red lines).
@@ 4.6 Stellar Populations
In this section we briefly consider the resolved stellar populations of the fields under study. Figure 8 shows color-magnitude diagrams of these galaxies with PARSEC isochrones (Bressan et al., 2012; Marigo et al., 2013) of various ages and metallicities overplotted. The CMDs demonstrate the full A1+A2 fields, corresponding to outer regions of each galaxy $\sim$20 kpc from their centers. The central galaxy of the Fornax cluster is the giant elliptical NGC 1399. Spectral observations of NGC 1399 made with MUSE (Vaughan et al., 2018) revealed a smooth decrease of the metallicity [Fe/H] from the center to the periphery, reaching a value slightly less than solar at the effective radius ($r_{e}$). To better understand the ages and metallicities of the resolved stars of NGC 1399, we take the PARSEC isochrones in the JWST/NIRCam system, using the distances derived in this work (Table 4). In general, the CMDs agree best with stellar population models with metallicity $Z$=0.008, somewhat less than solar metallicity ($Z_{\odot}$=0.015). Blue isochrones correspond to the E-AGB and TP-AGB evolutionary stages of 1 Gyr old stars. Thus, even in the outer regions of this galaxy, we can expect a number of stars with relatively young ages (about 1–2 Gyrs). Older red giant branch (RGB) and AGB stars, up to the oldest at 13.7 Gyr, appear to form the bulk of the population. A fairly well populated “tail” stretching to the right and downwards is clearly visible, showing the presence of red giants and AGB stars of solar metallicity and higher.
@@ 5 The Next Step: Calibrating the SBF Distance Scale
The SBF technique (Tonry & Schneider, 1988) is a statistical measurement of the brightness of the luminosity-weighted mean luminosity of a population of stars, which is dominated by RGB stars. Nearer galaxies show a mottled appearance that fades into smoothness for more distant galaxies. Unlike the TRGB, SBF measurements do not require resolution of individual RGB stars, and can be measured much further away, to distances beyond 100 Mpc with HST (Blakeslee et al., 2021a), and potentially much further with JWST (Cantiello & Blakeslee, 2023). The typical internal SBF distance errors (not including systematic uncertainty) with HST are less than 5% per galaxy, putting SBF on the same statistical level as type Ia supernovae, and capable of reaching well into the Hubble flow. It is therefore advantageous to have a dual measurement of SBF and TRGB for the same galaxies using the same filters and cameras to calibrate SBF, eliminating as many sources of systematic uncertainty as possible. This will allow us to directly address the persistent conflict between local distance ladder measurements of the Hubble constant $H_{0}$ and the value inferred from $\Lambda$CDM cosmological models (Planck Collaboration et al., 2020), sometimes referred to as the Hubble tension (Di Valentino et al., 2021). This paper outlines the first steps towards our goal of establishing a distance ladder with no reliance on Cepheids or supernovae, with the ultimate goal of reaching a distance precision approaching 1%.

### v2

@@ 1 Introduction
For at least a half a century, the Virgo Cluster has played an important role in the establishment of the extragalactic distance scale (Sandage & Tammann, 1976). However, the depth of the virializing region of this cluster, about ±$\pm 10\%$ of its distance, is an appreciated hindrance (increasing the scatter in the mean distance to the cluster), made more difficult by the ongoing accretion of new members (Tully & Shaya, 1984; Cantiello et al., 2024). The Fornax Cluster has played a more limited role, but it is more compact with a depth about ±$\pm 5\%$ of its distance and little evidence of recent accretion. Fornax is only slightly more distant than Virgo; surface brightness fluctuation (SBF) measurements for 43 galaxies in Fornax and 91 galaxies in the Virgo vicinity using the Advanced Camera for Surveys (ACS) on Hubble Space Telescope (HST), imply a Fornax/Virgo distance ratio $d_{F}/d_{V}=1.21\pm 0.02$ (Blakeslee et al., 2009).
@@
[…]
@@ 2 Data
The three galaxies presented in this work are the first of 14 targets of Cycle 2 JWST program GO-3055 (Tully et al., 2023a) to link the TRGB and SBF scales through simultaneous observations tailored to each method. The split fields of the JWST NIRCam detector (Gardner et al., 2023; Rieke et al., 2023) are ideal for the experiment. One field covers the high surface brightness core of a galaxy for the SBF measurement, while the other field extends to the galactic halo where the TRGB measurement can be made with minimal confusion from crowding. SBF can also be measured in the peripheral fields, providing a direct overlap with TRGB. The primary observations are made in the $F090W$ and $F150W$ filters. $F090W$ is optimal for TRGB and SBF because of the minimal variance of RGB magnitudes at lower metallicities (Anand et al., 2024). $F150W$ is another good choice for SBF because the fluctuation signal is very bright at this wavelength for the old, metal-rich populations of elliptical galaxies, making SBF measurable to distances beyond 100 Mpc (Jensen et al., 2015). Simultaneous long-wavelength observations are made in both $F277W$ and $F356W$, although we do not make use of these filters in this work. The integration times are set by the requirements of the TRGB observations. The SBF signal is abundantly strong in the same exposure time. The deep multi-wavelength imaging of the selected galaxies is also useful for many other ancillary science cases.
@@
2023-10-12 NGC 1399 o002_t002 F090W 7730.474
2023-10-12 NGC 1399 o002_t002 F150W 1932.619
2023-11-09 NGC 1380 o001_t001 F090W 7730.474
2023-11-09 NGC 1380 o001_t001 F150W 1932.619
2023-11-14 NGC 1404 o003_t003 F090W 7730.474
2023-11-14 NGC 1404 o003_t003 F150W 1932.619
Table 1: Observation log for the data used in this paper. All the images were processed with jwst_1140.pmap or jwst_1147.pmap context versions, which include the NIRCam flat-field and flux calibration updates introduced in jwst_1125.pmap and jwst_1126.pmap, respectively. The name of all observation IDs start with “jw03055-” and end with “_nircam_clear-f090w” or “_nircam_clear-f150w”. The times given are the actual measurement times (TMEASURE), and not the total exposure times (EFFEXPTM), which include some overhead.
Table 1 provides a summary of the data we use in this paper. All the data was reduced with context versions [^1] See https://jwst-crds.stsci.edu/display_all_contexts/ for more information on each version. 1140.pmap or 1147.pmap (with no meaningful changes for NIRCam imaging between the two), and contain all reference file updates as of January 2024, including the latest updates to the photometric zeropoints, provided in 1126.pmap. The observational setup with NIRCam was identical for each of the three galaxies presented in this paper. For each $F090W$ observation, we used a MEDIUM8 readout pattern with 7 groups/integration, 3 integrations/exposure, and four subpixel dithers. For each $F150W$ observation, we used a SHALLOW4 readout pattern, with 10 groups/integration, 1 integration/exposure, and four subpixel dithers.
Footprints of our NIRCam observations are shown in Fig. 1, overlaid on a one square degree image from the Digitized Sky Survey. From this figure, the extended spatial NIRCam coverage is clear. To further highlight the dynamic range probed by NIRCam in a single pointing, we show color ($F090W+F150W+F356W$) images of individual chips from the NGC 1380 visit in Fig. 2. The core of this galaxy is centered on the B4 chip; a small dust lane in the center is readily apparent. While individual stars cannot be resolved in the inner regions due to crowding, a few globular clusters are visible scattered throughout as marginally resolved sources. Some background galaxies can also be seen. The two chips furthest from the core of the galaxy are A1 and A2, and we show a color image from the former. There are thousands of giant branch stars which populate this frame. Importantly, they can be easily separated from one another, due to the significantly lower surface brightness at this radius. A diverse population of background galaxies is also visible.
@@ 3 Data Reduction
For our photometry, we use the recommended parameters provided in the recent work of Weisz et al. (2024). We perform our PSF photometry directly on the Stage 2 cal images, while using a Stage 3 i2d image from the JWST pipeline as the “reference frame” for mutual image alignment. We choose $F090W$ for the reference frame, as it has the highest angular resolution, and our observations are also the deepest in that filter. We filter our final object list based on the quality criteria presented in Anand et al. (2024), which are based on the recommendations provided by the ERS program team in Warfield et al. (2023). We only include sources which meet the following DOLPHOT criteria in both filters in our final object list (except for type, which is not a filter specific output): (1) Crowding $<$ 0.5; (2) $\mathrm{(Sharpness)^{2}}$leq 0.01; (3) Object Type $\leq$ 2; (4) S/N $\geq$ 5; (5) Error Flag $\leq$ 2. All of our PSF photometry is transformed to the Vega system (and not the Sirius–Vega variant [^2] By default, DOLPHOT uses the Sirius–Vega system. This can be switched to the Vega system by switching the filters_vega.dat $\rightarrow$ filters.dat in DOLPHOT’s NIRCam installation folder. The difference between the two photometric systems is $\sim$0.04 mag in $F090W$, and $\sim$0.02 mag in $F150W$.).
We opt to utilize the “–etctime” option in the DOLPHOT/NIRCam module, which rectifies the effective exposure time provided in the JWST image headers (EFFEXPTM) to use the appropriate measurement time (TMEASURE) for derivation of the underlying S/N. The use of this option does not significantly affect the flux measurements themselves (due to the underlying image units of surface brightness); instead it aims to provide more realistic S/N estimates for each individual source (Weisz et al., 2024).
@@
Color-magnitude diagrams (CMDs) for each target can be seen in Fig. 3. These CMDs are limited to the two SW NIRCam chips that are furthest from the galaxy’s core (A1+A2). For NGC 1399, we further restrict the points to just those that are in the furthest quarter of each chip, for visual clarity. In each CMD, the vast majority of resolved stars are those on the red-giant branch (RGB), with the TRGB visible at $m_{F090W}\simeq 27$ mag ($D=19$ Mpc) for each target. The magnitude limit of the CMD for NGC 1399 is $\sim 0.3$ mag brighter than for the other two targets, likely due to crowding effects at the faint end (NGC 1399 is the cD galaxy in the Fornax Cluster, and thus more extended). It can also be seen that the RGB contains stars of a vast range of metallicities; for our TRGB work, we will only use the lower metallicity stars, which have little-to-no variation in their magnitude when observed in $F090W$ (see the discussion in §2 of Anand et al. 2024). A further description of the underlying stellar populations can be found in §4.6.
@@ 4 TRGB Measurements
For each target, we perform our measurements on each of the two chips that are furthest from the galaxy’s core. Due to the observational setup for our program, these are the A1 and A2 chips for all three targets. We analyze the data for both chips separately, reducing any concerns regarding the chip-to-chip zeropoint offsets. We note that for most standard stars, these offsets are at the level of 0.00–0.03 mag [^3] See the details provided at https://jwst-docs.stsci.edu/jwst-data-calibration-considerations/jwst-calibration-uncertainties for more information regarding calibration uncertainties.. For NGC 1399, we also limit our analysis to the least crowded quarter of each of the two chips (via a simple spatial cut on the x-axis) in an attempt to mitigate the effects of the (modest) crowding present (which is more of an issue for the edge-detection routine, as it does not employ the use of artificial stars).
@@
Figure 4: Color-magnitude diagrams and edge-detection measurements for each of our six fields in three galaxies. The maximum-likelihood TRGB results are also plotted for comparison, but are nearly indistinguishable from the edge-detection results except for the NGC 1399 fields (which suffer from a modest amount of photometric bias). The grey-shaded region indicates the color range used for the TRGB measurements.
@@ 4.4 Distances
To provide an absolute scaling for our TRGB measurements, we utilize the JWST/NIRCam TRGB calibration provided by Anand et al. (2024). They use data taken in the outer regions of the megamaser host NGC 4258 to provide a direct calibration in our filter of interest ($F090W$). The baseline absolute calibration they provide is $M_{\mathrm{TRGB}}^{F090W}=-4.362\pm 0.033$ (stat) ±$\pm 0.045$ (sys) mag. They also note that this baseline calibration may be further tailored to individual measurement techniques: a value of $M_{\mathrm{TRGB}}^{F090W}=-4.347$ mag is more appropriate for techniques which employ the use of LF fitting with artificial stars (e.g. Makarov et al. 2006) and take into account photometric bias, errors, and incompleteness. It has been shown that the model-LF fitting techniques can provide measurements that are up to $\sim$0.05 magnitudes fainter than their edge-detection counterparts (Anand et al., 2022), consistent with the 0.03 mag offset in their absolute JWST calibration (Anand et al., 2024). We also note that several modest changes in the recommended DOLPHOT parameters may result in minor offsets in the final photometry, though Weisz et al. (2024) note that the performance is comparable. [^5] An updated NIRCam TRGB calibration with the new DOLPHOT parameters and PSF models, along with a NIRISS TRGB calibration will be provided in a future, separate work (Anand et al. in prep). Preliminary results from that work show that the net effect is insignificant ($<$0.01 mag). To account for any potential mismatches, we choose to include an additional 0.03 mag contribution to the overall systematic error term. The magnitude of this term is also provided by Anand et al. (2024), who obtain this value from several other works in the literature which examine the systematic effects of changing DOLPHOT parameters on the output photometry (Jang et al., 2021; Jang, 2023). Our adopted absolute calibration now becomes $M_{\mathrm{TRGB}}^{F090W}=-4.347\pm 0.033$ (stat) ±$\pm 0.054$ (sys) mag (see Table 3 in Anand et al. 2024 for all of the included error terms).
@@
Figure 7: Comparison of two CMDs (down to S/N = 5) for NGC 1404 used to measure the TRGB distance to this galaxy. The left panel shows the reduction of the 34 orbit (crowding-limited) HST dataset, whereas the right panel shows the reduction of the 4.74 hour JWST dataset. The y-axes of the two panels are set to match at the measured levels of the TRGB (denoted with dashed red lines). Insets show the respective observing footprints overlaid on ${}^{\prime\prime}\times$12′′ DSS imaging, with the colored regions denoting the regions shown in the CMDs and used for the TRGB measurements.
@@ 4.6 Stellar Populations
In this section we briefly consider the resolved stellar populations of the fields under study. Figure 8 shows color-magnitude diagrams of these galaxies with PARSEC isochrones (Bressan et al., 2012; Marigo et al., 2013) of various ages and metallicities overplotted. The CMDs demonstrate the full A1+A2 fields, corresponding to outer regions of each galaxy $\sim$20 kpc from their centers. The central galaxy of the Fornax cluster is the giant elliptical NGC 1399. Spectral observations of NGC 1399 made with MUSE (Vaughan et al., 2018) revealed a smooth decrease of the metallicity [Fe/H] from the center to the periphery, reaching a value slightly less than solar at the effective radius ($r_{e}$). To better understand the ages and metallicities of the resolved stars of NGC 1399, we take the PARSEC isochrones in the JWST/NIRCam system, using the distances derived in this work (Table 4). In general, the CMDs agree best with stellar population models with metallicity $Z$=0.008, somewhat less than solar metallicity ($Z_{\odot}$=0.015). Blue isochrones correspond to the E-AGB and TP-AGB evolutionary stages of 1 Gyr old stars. Thus, even in the outer regions of this galaxy, we can expect a number of stars with relatively young ages (about 1–2 Gyrs). Older RGB and AGB stars, up to the oldest at 13.7 Gyr, appear to form the bulk of the population. A fairly well populated “tail” stretching to the right and downwards is clearly visible, showing the presence of red giants and AGB stars of solar metallicity and higher.
@@ 5 The Next Step: Calibrating the SBF Distance Scale
The SBF technique (Tonry & Schneider, 1988) is a statistical measurement of the spatial fluctuations in surface brightness of a population of stars, which is dominated by the luminous RGB stars in an old smooth elliptical galaxy. Nearer galaxies show a mottled appearance that fades into smoothness for more distant galaxies. Unlike the TRGB, SBF measurements do not require resolution of individual RGB stars, and can be measured much further away, to distances beyond 100 Mpc with HST (Blakeslee et al., 2021a), and potentially much further with JWST (Cantiello & Blakeslee, 2023). The typical internal SBF distance errors (not including systematic uncertainty) with HST are less than 5% per galaxy, putting SBF on the same statistical level as type Ia supernovae, and capable of reaching well into the Hubble flow. It is therefore advantageous to have a dual measurement of SBF and TRGB for the same galaxies using the same filters and cameras to calibrate SBF, eliminating as many sources of systematic uncertainty as possible. This will allow us to directly address the persistent conflict between local distance ladder measurements of the Hubble constant $H_{0}$ and the value inferred from $\Lambda$CDM cosmological models (Planck Collaboration et al., 2020), sometimes referred to as the Hubble tension (Di Valentino et al., 2021). This paper outlines the first steps towards our goal of establishing a distance ladder with no reliance on Cepheids or supernovae, with the ultimate goal of reaching a distance precision approaching 1%.